# Supporting

### v1

@@ 3.1 Result for continuous outcomes
For continuous outcomes under Assumptions 11118888, assume the following conditions hold as $n\to\infty$: $\lambda_{n}\to 0$, $h_{n}\to 0$, $\lambda_{n}n^{\min\{2\gamma,1\}}\to\infty$, and $\lambda_{n}nh_{n}^{2}\to\infty$. Suppose further that either $\pi^{0}=\pi^{*}$ or $g^{0}=g^{*}$ holds, and the initial interval $(l^{1},r^{1})$ for Algorithm 1 contains the optimal quantile $q^{*}$. Let $\zeta>0$, $20<v<$ 2, $\sigma_{n}=c_{1}\lambda_{n}^{-1/\{(\beta+1)d\}}$, $\epsilon_{n}=c_{2}n^{-1/2}$, and $\kappa_{n}=c_{3}n^{-1/2}$, where $c_{1},c_{2},c_{3}>0$ are constants. Then we have
@@
Since the quantile OTRs may be non-unique Zhou et al., (2023), Theorem 3.1 primarily focuses on the quantile value consistency, rather than the consistency of the parameters for the quantile OTRs themselves. Theorem 3.1 reveals that the estimated quantile OTR $\hat{d}$ of our method is doubly robust in the sense that the quantile given $\hat{d}$ will be consistent to the optimal quantile if either the model for the propensity score $\pi^{*}$ or the conditional survival function $g^{*}$ is correctly specified, i.e., either $\pi^{0}=\pi^{*}$ or $g^{0}=g^{*}$.
@@
The optimal rate for the value function is $\max\{n^{-2\beta\eta/[\{6\beta+2+(2-v)(1+\zeta)\}(\eta+1)]},n^{-2\gamma\beta\eta/\{(3\beta+1)(\eta+1)\}}\}$ up to a logarithm with $h_{n}=1/\log n$ and $\lambda_{n}=\max\{n^{-2(\beta+1)/\{6\beta+2+(2-v)(1+\zeta)\}},n^{-2\gamma(\beta+1)/(3\beta+1)}\}$. The rate consists of two parts: the first part reflects the rate of convergence in estimating the quantile OTRs; the second part is related to nuisance function estimation. When data are well separated, which implies that $\beta$ and $\eta$ in Assumptions 6 and 7 can be arbitrarily large, then the convergence rate almost achieves the rate $\max\{n^{-1/3},n^{-2\gamma/3}\}$. When the nuisance models are estimated with convergence rates $\Delta_{n}=O_{p}(n^{-1/2})$ such that $\gamma=1/2$, which is satisfied for most parametric estimators, the rate for the value function is about $n^{-1/3}$. Our result differs from those in Zhao et al., 2015b ; Zhou et al., (2017), as we focus on the convergence rate of the quantile difference, whereas the existing literature primarily investigates the difference in expectations, i.e., $E\{Y(\hat{d})\}-\max_{d}E\{Y(d)\}$.
@@
$n^{-2\gamma\beta/(3\beta+1)}\}$ as Section LABEL:s6.7 of the supplementary material suggests. When the nuisance functions are estimated with convergence rates $\Delta_{n}=O_{p}(n^{-1/2})$ and the data are well separated, the convergence rate for this oracle method is about $n^{-1/3}$. This indicates that the convergence rate of our method is close to the convergence rate of the oracle method in the well-separated case.
@@ 3.2 Result for discrete outcomes
In this subsection, we establish theoretical guarantees for our method when the outcomes are discrete. As our approach is designed to maximize the smoothed quantile, we begin by demonstrating the convergence rate of $Q^{m}_{\tau}\{Y(\hat{d})\}-Q^{m}_{\tau}\{Y(\tilde{d}^{*})\}$ when a Gaussian kernel is used to construct $\hat{d}$, where $Q_{\tau}^{m}$ is the smoothed quantile and $\tilde{d}^{*}=\operatorname*{arg\,max}_{d}Q_{\tau}^{m}\{Y(d)\}$. We then establish that $Q_{\tau}\{Y(\hat{d})\}$ equals the optimal quantile $\max_{d}Q_{\tau}\{Y({d})\}$ with probability tending to 1. Namely, $\lim_{n\to\infty}P[Q_{\tau}\{Y(\hat{d})\}=\max_{d}Q_{\tau}\{Y({d})\}]=1$, a special asymptotic result that is typically observed and of interest in the context of discrete quantile function Machado and Silva, (2005); Chen and Lazar, (2010). Finally, we show that our method is consistent in quantile value and effective under the homoscedastic treatment effect assumption.
@@
Assumption 9 holds for binary outcomes and count outcomes following the conditional Poisson distribution. See Section LABEL:s6.5 of the supplementary material for details.
Since the smoothing technique maps discrete outcomes to continuous ones, the proof strategy employed in Theorem 3.1 can be directly extended to establish the following convergence rate result for $Q^{m}_{\tau}\{Y(\hat{d})\}-Q^{m}_{\tau}\{Y(\tilde{d}^{*})\}$, with Assumptions 2 and 6–8 appropriately modified to Assumptions $2^{*}$ and $6^{*}$$8^{*}$. The modified assumptions are provided in Section LABEL:s1.2 of the supplementary material. Then, using the convergence of $Q^{m}_{\tau}\{Y(\hat{d})\}-Q^{m}_{\tau}\{Y(\tilde{d}^{*})\}$, we can establish that $Q_{\tau}\{Y(\hat{d})\}$ equals the optimal quantile $\max_{d}Q_{\tau}\{Y({d})\}$ with probability tending to 1. These results are summarized in Theorem 3.2.
@@
For discrete outcomes under Assumptions $1,2^{*}$, $3-5$, $6^{*}-8^{*}$, assume the following conditions hold as $n\to\infty$: $\lambda_{n}\to 0$, $h_{n}\to 0$, $\lambda_{n}n^{\min\{2\gamma,1\}}\to\infty$, and $\lambda_{n}nh_{n}^{2}\to\infty$. Suppose further that either $\pi^{0}=\pi^{*}$ or $g^{0}=g^{*}$ holds, and the initial interval $(l^{1},r^{1})$ contains the optimal smoothed quantile $\max_{d}Q_{\tau}^{m}\{Y(d)\}$. Let $\zeta>0$, $20<v<$ 2, $\sigma_{n}=c_{1}\lambda_{n}^{-1/\{(\beta+1)d\}}$, $\epsilon_{n}=c_{2}n^{-1/2}$, and $\kappa_{n}=c_{3}n^{-1/2}$, where $c_{1},c_{2},c_{3}>0$ are constants. Then we have
@@
The assumptions underpinning Theorem 3.2 closely mirror those required for Theorem 3.1. Consequently, we refer the reader to Section 3.1 for a more comprehensive discussion of these assumptions. The part (i) of this theorem shows that the convergence rate of $Q_{\tau}^{m}\{Y(\hat{d})\}-Q_{\tau}^{m}\{Y(\tilde{d}^{*})\}$ for a discrete outcome is the same as the convergence rate of $Q_{\tau}\{Y(\hat{d})\}-Q_{\tau}\{Y(d^{*})\}$ for a continuous outcome, as established in Theorem 3.1. Thus, a comprehensive discussion of the convergence rate can be found in Section 3.1.
The part (ii) of Theorem 3.2 demonstrates that $Q_{\tau}\{Y(\hat{d})\}=\max_{d}Q_{\tau}\{Y({d})\}$ with probability tending to 1. Specifically, when the outcome $Y(d)$ is discrete, the quantile $Q_{\tau}\{Y({d})\}=\sup\{q:S(q,{d})>1-\tau\}$ is also discrete. Therefore, if $\hat{d}$ is a reliable estimator of the quantile OTRs with respect to the quantile value, $Q_{\tau}\{Y(\hat{d})\}$ should equal $\max_{d}Q_{\tau}\{Y({d})\}$ with probability tending to 1. Otherwise, $Q_{\tau}\{Y(\hat{d})\}$ would deviate significantly from the optimal quantile.
@@
The techniques employed to establish the part (ii) of Theorem 3.2 and Corollary 3.1 can likewise be applied to other methods that incorporate our smoothing approach, thereby demonstrating that those methods are consistent in quantile value and effective by combining our smoothing technique. Indeed, as shown in Sections LABEL:s1.2 and LABEL:s1.4 of the supplementary material, if a method satisfies smoothed quantile value consistency, it ensures quantile value consistency. Additionally, if the data satisfies the homoscedastic treatment effect assumption and boundedness $|Y|<M_{y}$, the method satisfying smoothed quantile value consistency is also effective.
@@ 4 Simulation
Case 2 (Continuous outcome with nonlinear regime): The continuous outcome $Y$ is generated by $Y=2-0.5X_{1}+0.5X_{2}^{2}+A(X_{1}+X_{2})+\mathrm{exp}\{-2-0.5X_{1}-0.5X_{2}+A(2+X_{1}+X_{2})\}\epsilon$, where $\epsilon$ follows a standard normal distribution. In this case, the quantile OTRs are nonlinear. We plot the boundaries of quantile OTRs for $\tau=0.25$ and $\tau=0.50$ in Figure LABEL:spfig1 of the supplementary material.
Case 3 (Discrete outcome with nonlinear regime): The discrete outcome $Y$ follows the negative binomial distribution with number of success $A+2$ and success probability $\mathrm{expit}\{-2-0.5X_{1}+0.25X_{2}^{2}+A(1+X_{1}+X_{2}^{2})\}$. Figure LABEL:spfig1 in the supplementary material illustrates the boundaries for quantile OTRs, which suggests that the quantile OTRs are nonlinear.
@@
3. Wang’s Method Wang et al., (2018): A doubly robust approach that leverages both the propensity score and conditional quantile information to ensure robustness. The propensity score is estimated using logistic regression models, as outlined in Section 2.4, while the conditional quantile is estimated via quantile regression models incorporating the basis function set $(\boldsymbol{X},A,\boldsymbol{X}A)$. Since the method is not directly applicable to discrete outcomes, we implement our smoothing technique for the estimated quantile function and then apply the genetic algorithm to maximize the estimated smoothed quantile to derive the quantile OTRs. See Section LABEL:s7.2 of the supplementary material.
4. QIQ-Learning Linn et al., (2017): A quantile interactive Q-learning approach that estimates the conditional survival function $g^{*}(q;\boldsymbol{x},a)$ and transforms it into quantile OTRs. The conditional survival function is estimated according to the procedure detailed in Section 2.4. Since the method is not directly applicable to discrete outcomes, we implement our smoothing technique for the estimated conditional survival function and then transform the estimated conditional smoothed survival function into quantile OTRs. See Section LABEL:s7.2 of the supplementary material.
The simulations are conducted with a sample size of 500, repeated 200 times, and assessed using a large independent test set of 10,000 subjects. In this study, we utilize two performance indicators: the value function given the estimated quantile OTRs evaluated in the test set and the misclassification rate (MR) by comparing the true and the estimated quantile OTRs in test samples. For Cases 1 and 2, the value function is given by the quantile $Q_{\tau}$, while for Case 3, it is defined as the smoothed quantile $Q_{\tau}^{m}$. The results are reported in Table 1.
Table 1: Mean (with standard deviation in the parenthesis) of value functions and misclassification rates evaluated on independent test data for three cases. The best result in each scenario is highlighted in bold.
@@
Case 1 (optimal values: 0.485 and 2.789)
@@
Case 2 (optimal values: 2.250 and 2.783)
@@
Case 3 (optimal values: 3.349 and 8.853)
@@
Additional simulations incorporating more covariates, varying sample sizes, and alternative performance indicators, such as computational time, are presented in the supplementary material. Notably, our methods SCL-Linear and SCL-Gaussian are computationally more efficient than Wang’s method. To further validate the doubly robust property of our method in the presence of the model misspecification of the propensity score $\pi^{*}$ or the conditional survival function $g^{*}$ for both the linear kernel and the Gaussian kernel, we include simulation results in the supplementary material.
@@ 5 Real data analysis
We illustrate our method on the ACTG175 dataset from the R package speff2trial, which contains measurements on 2139 HIV-infected patients. The patients were randomized to four treatment arms: zidovudine (AZT) monotherapy, AZT+didanosine (ddI), AZT+zalcitabine (ddC), and ddI monotherapy. The outcome we consider is the CD4 cell count (per cubic millimeter) at 20 ± 5 weeks post the baseline. A larger value indicates a better outcome. Same as Wang et al., (2018), we treat the CD4 cell count as a continuous outcome. We focus on assigning treatments, specifically the AZT+ddI combination therapy or the ddI monotherapy, to patients, as Wang et al., (2018) revealed these two treatments exhibit heteroscedastic treatment effects.
We consider all twelve variables measured at baseline as the covariates for each subject followed by Fan et al., (2017). Five of the twelve covariates are continuous or count variables: age (years), weight (kilogram), Karnofsky score (on a scale of 0-100), CD4 cell counts at baseline (per cubic millimeter), and CD8 cell counts at baseline (per cubic millimeter). The rest seven are binary: hemophilia (0=no, 1=yes), homosexual activity (0=no, 1=yes), history of intravenous drug use (0=no, 1=yes), race (0=white, 1=non-white), gender (0=female, 1=male), antiretroviral history (0=naive, 1=experienced), and symptomatic indicator (0=asymptomatic, 1=symptomatic).
We consider $\tau=0.25,0.50$, and $0.75$. The performance of our method is compared with other approaches discussed in Section 4 using cross-validation. Specifically, we randomly split the dataset into equal training and testing sets 200 times. For each split, we estimate the quantile OTRs using each method based on the training set. We then evaluate the performance of the estimated quantile OTRs for each method by computing the estimated quantile $\widehat{Q}_{\tau}\{Y(\hat{d})\}$ using the testing set, where $\widehat{Q}_{\tau}\{Y(\hat{d})\}$ is calculated as in (2). The results are shown in Table 2. It is evident that our method, SCL-Gaussian, yields the largest quantile values at all the quantile levels.
Table 2: Mean (standard deviation) of value function from the cross-validation procedure using different methods. The best result in each scenario is highlighted in bold.

### v2

@@ 3.1 Result for continuous outcomes
For continuous outcomes under Assumptions 1188, assume the following conditions hold as $n\to\infty$: $\lambda_{n}\to 0$, $h_{n}\to 0$, $\lambda_{n}n^{\min\{2\gamma,1\}}\to\infty$, and $\lambda_{n}nh_{n}^{2}\to\infty$. Suppose further that the initial interval $(l^{1},r^{1})$ for Algorithm 1 contains the optimal quantile $q^{*}$. Let $\zeta>0$, $0<v<2$, $\sigma_{n}=c_{1}\lambda_{n}^{-1/\{(\beta+1)d\}}$, $\epsilon_{n}=c_{2}n^{-1/2}$, and $\kappa_{n}=c_{3}n^{-1/2}$, where $c_{1},c_{2},c_{3}>0$ are constants. Then we have
@@
Since the quantile OTRs may be non-unique zhouTransformationInvariantLearningOptimal2022, Theorem 3.1 primarily focuses on the quantile value consistency, rather than the consistency of the parameters for the quantile OTRs themselves. As Assumption 5 holds when either the model for the propensity score $\pi^{*}$ or the conditional survival function $g^{*}$ is correctly specified, i.e., either $\pi^{0}=\pi^{*}$ or $g^{0}=g^{*}$, Theorem 3.1 reveals that the estimated quantile OTR $\hat{d}$ of our method is doubly robust in the sense that the quantile given $\hat{d}$ will be consistent with the optimal quantile. The bound in Theorem 3.1 depends on the convergence rates of the estimated propensity score $\hat{\pi}$ and the estimated conditional survival function $\hat{g}$ toward their limits $\pi^{0}$ and $g^{0}$, as captured by $\Delta_{n}$ defined in Assumption 5. In general, the same bound for $\Delta_{n}$, and hence the same bound in Theorem 3.1, can be achieved even when only one of $\pi^{0}=\pi^{*}$ and $g^{0}=g^{*}$ holds. This is a distinctive asymptotic property for doubly robust classification, on which our method is built zhao2015doubly. We refer the reader to Section 5.5 of the supplementary material for further details.
@@
The optimal rate for the value function is $\max\{n^{-2\beta\eta/[\{6\beta+2+(2-v)(1+\zeta)\}(\eta+1)]},n^{-2\gamma\beta\eta/\{(3\beta+1)(\eta+1)\}}\}$ up to a logarithm with $h_{n}=1/\log n$ and $\lambda_{n}=\max\{n^{-2(\beta+1)/\{6\beta+2+(2-v)(1+\zeta)\}},n^{-2\gamma(\beta+1)/(3\beta+1)}\}$. The rate consists of two parts: the first part reflects the rate of convergence in estimating the quantile OTRs; the second part is related to nuisance function estimation. When data are well separated, which implies that $\beta$ and $\eta$ in Assumptions 6 and 7 can be arbitrarily large, then the convergence rate almost achieves the rate $\max\{n^{-1/3},n^{-2\gamma/3}\}$. When the nuisance models are estimated with convergence rates $\Delta_{n}=O_{p}(n^{-1/2})$ such that $\gamma=1/2$, which is satisfied for most parametric estimators, the rate for the value function is about $n^{-1/3}$. Our result differs from those in zhao2015doubly; zhou2017residual, as we focus on the convergence rate of the quantile difference, whereas the existing literature primarily investigates the difference in expectations, i.e., $E\{Y(\hat{d})\}-\max_{d}E\{Y(d)\}$.
@@
$n^{-2\gamma\beta/(3\beta+1)}\}$ as Section S5.6 of the supplementary material suggests. When the nuisance functions are estimated with convergence rates $\Delta_{n}=O_{p}(n^{-1/2})$ and the data are well separated, the convergence rate for this oracle method is about $n^{-1/3}$. This indicates that the convergence rate of our method is close to the convergence rate of the oracle method in the well-separated case.
@@ 3.2 Result for discrete outcomes
[…]
@@
Assumption 9 holds for binary outcomes and count outcomes following the conditional Poisson distribution. See Section S5.2 of the supplementary material for details.
[…]
@@
For discrete outcomes under Assumptions $1,2^{*}$, 353-5, $6^{*}-8^{*}$, assume the following conditions hold as $n\to\infty$: $\lambda_{n}\to 0$, $h_{n}\to 0$, $\lambda_{n}n^{\min\{2\gamma,1\}}\to\infty$, and $\lambda_{n}nh_{n}^{2}\to\infty$. Suppose further that the initial interval $(l^{1},r^{1})$ contains the optimal smoothed quantile $\max_{d}Q_{\tau}^{m}\{Y(d)\}$. Let $\zeta>0$, $0<v<2$, $\sigma_{n}=c_{1}\lambda_{n}^{-1/\{(\beta+1)d\}}$, $\epsilon_{n}=c_{2}n^{-1/2}$, and $\kappa_{n}=c_{3}n^{-1/2}$, where $c_{1},c_{2},c_{3}>0$ are constants. Then we have
@@
The assumptions underpinning Theorem 3.2 closely mirror those required for Theorem 3.1. Consequently, we refer the reader to Section 3.1 for a more comprehensive discussion of these assumptions. The part (i) of this theorem shows that the convergence rate $Q_{\tau}^{m}\{Y(\hat{d})\}-Q_{\tau}^{m}\{Y(\tilde{d}^{*})\} fo$r a discrete outcome is the same as the convergence rate of $Q_{\tau}\{Y(\hat{d})\}-Q_{\tau}\{Y(d^{*})\}$ for a continuous outcome, as established in Theorem 3.1. Thus, a comprehensive discussion of the convergence rate can be found in Section 3.1.
[…]
@@
The techniques employed to establish the part (ii) of Theorem 3.2 and Corollary 3.1 can likewise be applied to other methods that incorporate our smoothing approach, thereby demonstrating that those methods are consistent in quantile value and effective by combining our smoothing technique. Indeed, as shown in Sections S6.2 and S6.3 of the supplementary material, if a method satisfies smoothed quantile value consistency, it ensures quantile value consistency. Additionally, if the data satisfies the homoscedastic treatment effect assumption and boundedness $|Y|<M_{y}$, the method satisfying smoothed quantile value consistency is also effective.
@@ 4 Simulation
Case 2 (Continuous outcome with nonlinear regime): The continuous outcome $Y$ is generated by $Y=2-0.5X_{1}+0.5X_{2}^{2}+A(X_{1}+X_{2})+\mathrm{exp}\{-2-0.5X_{1}-0.5X_{2}+A(2+X_{1}+X_{2})\}\epsilon$, where $\epsilon$ follows a standard normal distribution. In this case, the quantile OTRs are nonlinear. We plot the boundaries of quantile OTRs for $\tau=0.25$ and $\tau=0.50$ in Figures S4 and S5 of the supplementary material.
Case 3 (Discrete outcome with nonlinear regime): The discrete outcome $Y$ follows the negative binomial distribution with number of success $2A+1$ and success probability $\mathrm{expit}\{-2.5-0.5X_{1}+0.25X_{2}^{2}+A(1+X_{1}+X_{2}^{2})\}$. Figures S6 and S7 in the supplementary material illustrate the boundaries for quantile OTRs, which suggests that the quantile OTRs are nonlinear.
@@
3. Wang’s method wang2018quantileoptimal: A doubly robust approach that leverages both the propensity score and conditional quantile information to ensure robustness. The propensity score is estimated using logistic regression models, as outlined in Section 2.4, while the conditional quantile is estimated via quantile regression models incorporating the basis function set $(\boldsymbol{X},A,\boldsymbol{X}A)$. Since the method is not directly applicable to discrete outcomes, we implement our smoothing technique for the estimated quantile function and then apply the genetic algorithm to maximize the estimated smoothed quantile to derive the quantile OTRs. See Section S2.2 of the supplementary material.
4. QIQ-Learning linn2017interactive: A quantile interactive Q-learning approach that estimates the conditional survival function $g^{*}(q;\boldsymbol{x},a)$ and transforms it into quantile OTRs. The conditional survival function is estimated according to the procedure detailed in Section 2.4. Since the method is not directly applicable to discrete outcomes, we implement our smoothing technique for the estimated conditional survival function and then transform the estimated conditional smoothed survival function into quantile OTRs. See Section S2.2 of the supplementary material.
The simulations are conducted with a sample size of 500, repeated 200 times, and assessed using a large independent test set of 100,000 subjects. In this study, we utilize two performance indicators: the value function given the estimated quantile OTRs and the misclassification rate (MR) by comparing the true and the estimated quantile OTRs in test samples. For Cases 1 and 2, the value function is given by the quantile $Q_{\tau}$, while for Case 3, it is defined as the smoothed quantile $Q_{\tau}^{m}$. The results are reported in Table 1.
Table 1: Mean (with standard deviation in the parenthesis) of value functions (“Value”) and misclassification rates (“MR”) evaluated on independent test data for three cases. The best result in each scenario is highlighted in bold.
@@
Case 1 (optimal values: 0.485 for $\tau=0.25$ and 2.789 for $\tau=0.50$)
@@
Case 2 (optimal values: 2.250 for $\tau=0.25$ and 2.783 for $\tau=0.50$)
@@
Case 3 (optimal values: 3.349 for $\tau=0.25$ and 8.853 for $\tau=0.50$)
@@
In the supplementary material Section S3, we visualize the relationship between the estimated and true quantile OTRs. We also present additional simulation studies that consider a larger number of covariates and a range of sample sizes. Additional results, including computational time, the bias of the estimated optimal value function, and the estimated value function evaluated under the estimated quantile OTRs, are reported therein. Notably, our proposed approaches, SCL-Linear and SCL-Gaussian, demonstrate substantially greater computational efficiency than Wang’s method. In addition, we include simulation results in the supplementary material to validate the doubly robust property of our method under misspecification of the propensity score $\pi^{*}$ or the conditional survival function $g^{*}$, for both the linear kernel and the Gaussian kernel.
@@ 5 Real data analysis
We illustrate our method on the ACTG175 dataset from the R package speff2trial, which contains measurements on 2139 HIV-infected patients. The patients were randomized to four treatment arms: zidovudine (AZT) monotherapy, AZT+didanosine (ddI), AZT+zalcitabine (ddC), and ddI monotherapy. The outcome we consider is the CD4 cell count (per cubic millimeter) at 20 ± 5 weeks post the baseline. A larger value indicates a better outcome. Same as wang2018quantileoptimal, we treat the CD4 cell count as a continuous outcome. We focus on assigning treatments, specifically the AZT+ddI combination therapy or the ddI monotherapy, to patients, as wang2018quantileoptimal revealed these two treatments exhibit heteroscedastic treatment effects.
We consider all twelve variables measured at baseline as the covariates for each subject followed by fan2017concordanceassisted. Five of the twelve covariates are continuous or count variables: age (years), weight (kilogram), Karnofsky score (on a scale of 0-100), CD4 cell counts at baseline (per cubic millimeter), and CD8 cell counts at baseline (per cubic millimeter). The rest seven are binary: hemophilia (0=no, 1=yes), homosexual activity (0=no, 1=yes), history of intravenous drug use (0=no, 1=yes), race (0=white, 1=non-white), gender (0=female, 1=male), antiretroviral history (0=naive, 1=experienced), and symptomatic indicator (0=asymptomatic, 1=symptomatic).
[…]
Table 2: Mean (standard deviation) of value functions (“Value”) for different methods. The best result in each scenario is highlighted in bold.